Grade 11Physical Chemistry

1021 molecules are removed from 280 mg carbon mono oxide . Calculate the no of moles of carbon mono oxide

Ritik
8 Years agoGrade 11

1 Answer

Khimraj
ApprovedApproved Tutor Answer8 Years ago
initial mole of CO = 0.280/28 = 0.01 mole
removed mole = 1021/(6.022*1023) = 0.001661
SO remaining mole of CO = 0.01 – 0.001661 = 0.00834 mole.
Hope it clears. If you like answer then please aprove it.